OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4>?

OpenStudy (anonymous):

i think it is 0+2kpi

OpenStudy (anonymous):

thats no possiable as one of the given solutions..

OpenStudy (amistre64):

theres a relationship between cos and dot products, do you recall it?

OpenStudy (anonymous):

yes.

OpenStudy (amistre64):

then the angle is simple to be found knowing:\[|u||v|cos~\alpha=u \cdot v\]

OpenStudy (amistre64):

solve for alpha :)

OpenStudy (anonymous):

i remember the formula how do i plug the numbers into the formula?

OpenStudy (amistre64):

you know how to do a dot product right? and to find lengths of vectors?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

those are the numbers you need then, what do you get for: u.v and the lengths of u and v ?

OpenStudy (anonymous):

38

OpenStudy (amistre64):

u = <6, -1>; sqrt(37) v = <7, -4>; sqrt(65) --------------- u.v 42+5 = 49 |u||v| = sqrt(37*65) .... what does 38 mean?

OpenStudy (amistre64):

\[\cos~\alpha=\frac{u.v}{|u||v|}\] \[\cos~\alpha=\frac{49}{\sqrt{37*65}}\] \[\alpha=cos^{-1}\frac{49}{\sqrt{37*65}}\]

OpenStudy (anonymous):

the 38 came from 42+(-4)

OpenStudy (amistre64):

lol ... i do tend to add when i need to multiply when doing that

OpenStudy (amistre64):

-1*-4 is +4, and i did a +5 :)

OpenStudy (anonymous):

okay thank you i got 20.28?

OpenStudy (amistre64):

i got something closer to 10